RESONANCE-ENHANCED PHOTOASSOCIATIVE FORMATION OF GROUND-STATE Rb_{2} AND SPEC-TROSCOPY OF MIXED-CHARACTER EXCITED STATES

H. K. PECHKIS, D. WANG, Y. HUANG, \underline{E. E. EYLER}, P. L. GOULD, AND W. C. STWALLEY, Physics Department, University of Connecticut, Storrs, CT 06269; C. P. KOCH, Institut für Theoretische Physik, Freie Universität Berlin, 14195 Berlin, Germany.

We describe experimental and theoretical studies of the effects of resonant electronic state coupling on the formation of ultracold ground-state^{85}Rb_{2}. The molecules are formed by photoassociation of ultracold atoms in a MOT into the 0_{u}^{+} state converging to the 5S + 5P_{1/2} limit, followed by radiative decay into high vibrational levels of the ground electronic state, X ^{1}Σ_{g}^{+}. The populations of these high-v ground-state levels are monitored by resonance-enhanced two-photon ionization (R2PI) through the 2 ^{1}Σ_{u}^{+} state. We find that the populations of vibrational levels v′′=112-116 are far larger than can be accounted for by the Franck-Condon factors for 0_{u}^{+} ← X ^{1}Σ_{g}^{+} transitions. Further, the total number of ground-state molecules formed by this process exhibits oscillatory behavior as the PA laser is tuned through a succession of 0_{u}^{+} state vibrational levels. Both of these effects are explained by a new calculation of transition amplitudes that includes the resonant character of the spin-orbit coupling between the two 0_{u}^{+} states converging to the 5P_{1/2} and 5P_{3/2} limits. The resulting enhancement of more deeply bound ground-state molecule formation will be useful for future experiments on ultracold molecules.

We also describe evidence from our R2PI spectra for extensive singlet-triplet mixing between excited states of Rb_{2} at intermediate internuclear separations, apparently also induced by spin-orbit interactions. In particular, the 3 ^{1}Σ_{g}^{+} and 1 ^{1}Δ_{g} states converging to 5s + 4d have been observed in excitation from the a ^{3}Σ_{u}^{+} state,^{a} and the 2 ^{3}Π_{u} state has been observed in excitation from the X ^{1}Σ_{g}^{+} state.

aJ. Lozeille, A. Fioretti, C. Gabbanini, Y. Huang, H. K. Pechkis, D. Wang, P. L. Gould, E. E. Eyler, W. C. Stwalley, M. Aymar, and O. Dulieu, Eur. Phys. J. D 39, 261 (2006).